OpenStudy (anonymous):

Integrate x^4 * sqrt(1+x^2) dx

OpenStudy (anonymous):

I'll give a medal if it was solved

OpenStudy (b87lar):

Here is the solution but you may be disappointed it's not a beauty: \[\int x^4 \sqrt{1+x^2} dx = 1/48 (x \sqrt{x^2+1} (8 x^4+2 x^2-3)+3 \sinh^{-1}(x))+C\] To get this, start with substitution \[u=\tan x\]which will the lead you through several pages of calculations with trig integrals.

OpenStudy (b87lar):

sorry i meant to say substitution \[x=\tan u\]